# Search for Long-lived Particles at Future Lepton Colliders Using Deep Learning Techniques

### v2

@@ II Method
Event graphs are then input into a GNN-based heterogeneous architecture, as illustrated in Fig. 4. Features of different types of nodes and edges are first embedded into a high-dimension latent space and then forwarded to a Heterogeneous Detector Information Block (HDIB). The HDIB consists of two Detector Information Blocks (DIBs) [38][39] and two multilayer perceptrons (MLPs). The HDIB adopts a parameter sharing design between the tracker and the calorimeter types, specifically tailored for integrating information from different types of detectors. After the total $L$ layers of HDIBs, the final node embeddings of the tracker and the calorimeter are aggregated and forwarded to the decode layer to form the classification scores.
@@
The GNN training is performed with 8 GPUs and the batch size is 256 per GPU. The initial learning rate is $10^{-4}$, a dropout rate of 0.1 is used to avoid overfitting. The cross-entropy loss is minimized using the Adam optimizer without weight decay. The GNN model is trained with 30 epochs, and the network is validated at the end of each epoch. The model demonstrating the minimum validation loss value on the validation dataset is then applied to the final test dataset. The distributions of the training and testing losses for GNN are shown in Sec. 1.3 in ”Supplementary Information”.
@@
$R_{i}$ $\sqrt{\eta^{2}+\phi^{2}}$
tracker type edge between node $i$ and $j$ $|r_{i}-r_{j}|$, $r_{i}r_{j}$, $\eta_{i}-\eta_{j}$, $\phi_{i}-\phi_{j}$, $R_{i}-R_{j}$
@@
Systematic Uncertainties
Several sources of systematic uncertainties have been carefully evaluated:
  1. 1.
    Uncertainty from Higgs Boson Count: The uncertainty originating from the total number of Higgs bosons is estimated to be $1.0\%$ [40]. This accounts for variations in the production rate and experimental measurement of Higgs boson events.
  2. 2.
    Neural Network Training Variability: To assess the robustness of our machine learning model, we trained 50 CNNs with different initial seeds. The training uncertainty for the neural networks is estimated as half of the difference between the maximum and minimum efficiencies observed, which amounts to approximately $1.7\%$.
A quadratic sum of the above two uncertainties yields a total systematic uncertainty of $2.0\%$.
III Results
To fully explore the potential of searching for LLPs with lepton colliders, we develop a novel approach that directly utilizes the low-level detector information based on the energy, position, and time of detector hits, in contrast to the traditional approach which performs a full event reconstruction. Detailed information on the general analysis strategy and the deep learning networks used in this work is described in the ”Methods” section and ”Supplementary Information”. Fig. 2 shows the structure of the Convolutional Neural Networks model [22], and Fig. 4 illustrates the architecture of the Graph Neural Networks [23]. Both CNN and GNN methods have been evaluated for the efficiency of LLPs signals (type I and II signals combined) while keeping background free, the efficiency results are summarized in Table 2. The signal efficiencies of the CNN method range from 74% to 99%, while the GNN method efficiencies range from 79% to 97% for various LLP masses and lifetimes. Both efficiency numbers are similar. We calculate the signal acceptance based on CEPC detector geometry settings with the full detector coverage [24]. The signal acceptance factors are summarized in Table 3. We then multiply the signal efficiency by the signal acceptance to obtain the signal yield. The signal yield is shown in Figure 5.
Table 2: Signal efficiencies for different LLPs masses and lifetimes obtained with CNN-based and GNN-based approaches. All efficiency values are presented with their corresponding statistical uncertainties.
@@ III Results
Figure 5: Efficiency and acceptance of LLP Detection. The product of geometry acceptance and selection efficiency for LLPs with varying masses and lifetimes. Solid symbols represent data from a CNN-based approach, while hollow symbols are from a GNN-based approach. The varied shapes correspond to different assumed LLP masses, with circles for 50 GeV, diamonds for 10 GeV, and crosses for 1 GeV. The error bars representing statistical uncertainties have been omitted from the figure as their values are too small for effective visual representation.
To estimate the sensitivity of the study, we assume a null hypothesis for LLPs signals and obtain 95% Confidence Level upper limits [41] on the branching ratio $\mathcal{B}(H\rightarrow\text{LLPs})$ for the process $e^{+}e^{-}\to ZH(Z\to\text{inclusive},H\to X_{1}+X_{2}$). The analyzed samples have a 20 $20\text{~{}ab}^{-1}$ luminosity and about $4\times 10^{6}$ Higgs bosons. For the two LLPs signal types, we have considered the following scenarios:
@@
  • Type I and Type II signal yields have a fixed ratio. We define a parameter $\epsilon_{V}:=\frac{BR(X\to\nu\bar{\nu})}{BR(X\to q\bar{q})}$ as the ratio and set it with a fixed value of 0.2. A one-dimensional 95% Confidence Level upper limit on $\mathcal{B}(H\rightarrow\text{LLPs})$ is derived and shown in Figure 6a).
@@
  • Type I and Type II signal yields have a floating ratio $\epsilon_{V}$ with an allowed range of $10^{-6}$ and $100$. A one-dimensional 95% Confidence Level upper limit on $\mathcal{B}(H\rightarrow\text{LLPs})$ is derived and shown in Figure 6b).
@@
  • Type I and Type II signal yields has a floating ratio and two-dimensional 95% Confidence Level upper limits on $\mathcal{B}_{\text{2-jet}}$ and $\mathcal{B}_{\text{4-jet}}$ are derived. A bivariate statistical fit is performed to derive the upper limits and results are shown in Figure 7. More detailed results on 2-D upper limits can be found in Sec. 1.4 in ”Supplementary Information”.
@@
Both 1-D and 2-D exclusion limits are summarized in Table 4. As the results indicate, statistical uncertainties dominate in this analysis. For systematic uncertainties, refer to the ’Systematic Uncertainty’ section.
@@
Figure 6: One-dimensional constraints on Higgs boson decay to LLPs. 95% C.L. upper limit on the branching ratio (BR) for the Higgs boson ($H$) decay into pairs of LLPs ($X_{1}X_{2}$) via $e^{+}e^{-}\to ZH$, where $\epsilon_{V}$ is the ratio $\frac{BR(X\to\nu\bar{\nu})}{BR(X\to q\bar{q})}$. a): a fixed ratio $\epsilon_{V}=0.2$, b): a floating $\epsilon_{V}$. The shaded areas indicate statistical and systematic uncertainties combined.
@@
Figure 7: Two-dimensional constraints on Higgs Boson decay to LLPs. The 95% C.L. 2-D upper limit on ($\mathcal{B}_{\textrm{2-jet}}$, $\mathcal{B}_{\textrm{4-jet}}$) for the process $e^{+}e^{-}\to ZH$ for three LLPs masses 50 GeV (left), 10 GeV (middle), 1 GeV (right). Different colored lines indicate different LLPs lifetimes. The uncertainties on the limits are omitted and a few limits are scaled by a factor for better visibility.
@@
Table 4: The 95% C.L. exclusion limit on BR($h\to X_{1}X_{2}$) for all signal channels with both fixed and floating $\epsilon_{V}$. The limits include ±$\pm 1\sigma$ uncertainties after taking into account both statistical and systematic contributions.
@@
Fixed 1 $1.9^{+0.5}_{-0.3}$ $1.5^{+0.4}_{-0.2}$ $5.0^{+1.4}_{-0.7}$ $32.4^{+8.0}_{-3.6}$ $933.1^{+237.3}_{-74.4}$
10 $2.7^{+0.8}_{-0.2}$ $1.9^{+0.6}_{-0.2}$ $2.2^{+0.7}_{-0.3}$ $4.7^{+1.6}_{-0.6}$ $27.6^{+6.0}_{-2.0}$
50 $1.5^{+0.6}_{-0.1}$ $1.2^{+0.4}_{-0.1}$ $1.4^{+0.3}_{-0.2}$ $1.4^{+0.3}_{-0.2}$ $3.4^{+1.0}_{-0.3}$
Floating 1 $2.7^{+1.1}_{-1.8}$ $1.8^{+1.4}_{-1.0}$ $2.9^{+6.1}_{-0.4}$ $16.4^{+31.0}_{-0.8}$ $493.4^{+1090.1}_{-39.7}$
10 $3.8^{+0.3}_{-2.5}$ $4.3^{+0.5}_{-3.3}$ $1.3^{+2.7}_{-0.2}$ $2.8^{+5.8}_{-0.4}$ $15.3^{+24.6}_{-2.0}$
50 $2.8^{+1.0}_{-2.0}$ $2.4^{+0.7}_{-1.8}$ $1.5^{+1.3}_{-0.8}$ $1.5^{+1.3}_{-0.8}$ $2.1^{+4.2}_{-0.2}$
Discussion
We evaluate the effectiveness of our ML-based approach for LLPs detection at future lepton colliders by comparing our results with results at hadron colliders with the ATLAS experiment [13] and CMS experiment [14] as well as future HL-LHC experiments [25]. The comparison is based on four primary metrics: signal acceptance, selection efficiency, analysis strategy and signal yields:
@@ Discussion
  • Selection Efficiency: For hadron colliders, LLPs events typically trigger on displaced decays and/or large missing transverse energy. The LLPs trigger efficiency at the ATLAS experiment is estimated to be between $10^{-3}$ and $0.3$ [13]. Besides trigger efficiency, there are additional efficiencies involved such as displaced vertex/object reconstruction efficiencies which are typically in the order of a few percent. In contrast, LLPs event selection at lepton colliders can adapt to a triggerless approach owing to the clean environment. ML-based approach can be applied directly with low-level detector information without any event-level reconstruction. As a result, our ML-based approach with lepton colliders can achieve an overall selection efficiency as high as $99\%$, an improvement of several orders in magnitude when comparing with LHC or HL-LHC efficiencies.
@@
  • Analysis Strategy: Traditionally, analyses of LLPs conducted elsewhere have employed a selection-based method, which involves categorizing events into multiple subsets with different decay modes and orthogonal signal types. These analyses necessitate manual re-tuning and re-optimization for each subset and different LLPs mass and lifetime configurations. In contrast, our ML-based approach eliminates the need for manual categorization and optimization. Deep neural networks can be retrained with similar setup for each LLPs mass and lifetime, resulting in higher efficiencies compared to the selection-based method. Detailed comparison between two methods is shown in Sec. 1.1 in ”Supplementary Information”.
@@
  • Signal Yields and Upper Limits Comparison: LHC and HL-LHC can produce a significantly larger number of Higgs bosons compared to lepton colliders. Despite this, higher signal acceptance and selection efficiencies in our ML-based approach compensate for the relatively low number of Higgs bosons. We achieve upper limits as low as $1.2\times 10^{-6}$ on $\mathcal{B}(H\rightarrow\text{LLPs})$ with $4\times 10^{6}$ Higgs bosons. This upper limit is approximately three orders of magnitude better than the $10^{-3}$ limit observed at the ATLAS and CMS experiments with $10^{7}$ Higgs bosons and it is comparable to the projected HL-LHC limit with about $10^{8}$ expected Higgs bosons.
@@
Besides comparing with hadron colliders, we have also compared our result with a preliminary study [26] on the ILC [27] sensitivity with a traditional selection-based method. The ILC sensitivity study searches for long-lived dark photons produced in Higgstrahlung events via the Higgs portal. We have compared our result with the hadronic decay dark photo result since the event signature is similar. We have seen that the signal acceptance factors are similar between ILC and CEPC detectors but the signal efficiencies differ significantly. The signal efficiencies in the ILC study range from $0.1\%$ to $10\%$, which is at least an order of magnitude lower than ours. The upper limits on $\mathcal{B}(H\rightarrow\text{LLPs})$ in the ILC study are derived under the assumption of $100\%$ truth-level signal efficiency. Under this assumption, two results show similar sensitivities only in the low lifetime region ($<1~{}ns$) of LLPs. In the long lifetime region, for example, for a 1 GeV LLP with a lifetime of a few nanoseconds, our result yields an upper limit of about $5\times 10^{-6}$ which is an order of magnitude better than the ILC’s upper limit of $10^{-4}$.
In summary, we have developed an ML-based approach for LLP searches that outperforms traditional selection-based methods on almost all fronts. Moreover, our ML-based approach can be easily applied to other future lepton collider experiments, including the ILC, FCC [28], and CLIC [29].
There has been recent progress in LLP searches using ML-assisted approaches at hadron colliders [30, 31]. These results utilize ML techniques to perform object-level tagging looking for LLP signatures after full event reconstruction. The tagging performance relies on a limited number of high-level input variables. In our approach, both low-level information (based on hits) and global features of the event have been taken into account. Due to the harsh environment at hadron colliders, controlling the background becomes more challenging, which may require higher granularity and additional input features when constructing the neural networks. Additionally, advanced triggers with fast neural networks are necessary to efficiently capture LLP-like events at hadron colliders [32, 33].
IV Acknowledgement

### v5

@@ 3.3 Graph Neural Network
Event graphs are then input into a GNN-based heterogeneous architecture, as illustrated in Figure 5. Features of different types of nodes and edges are first embedded into a high-dimension latent space and then forwarded to a Heterogeneous Detector Information Block (HDIB). The HDIB consists of two Detector Information Blocks (DIBs) [47, 48]and two multilayer perceptrons (MLPs). The HDIB adopts a parameter-sharing design between the tracker and the calorimeter types, specifically tailored for integrating information from different types of detectors. After the total $L$ layers of HDIBs, the final node embeddings of the tracker and the calorimeter are aggregated and forwarded to the decoding layer to form the classification scores.
@@
The GNN training is performed with 8 GPUs and the batch size is 256 per GPU. The initial learning rate is $10^{-4}$, and a dropout rate of 0.1 is used to avoid overfitting. The cross-entropy loss is minimized using the Adam optimizer without weight decay. The GNN model is trained with 30 epochs, and the network is validated at the end of each epoch. The model demonstrating the minimum validation loss value on the validation dataset is then applied to the final test dataset. The distributions of the training and testing losses for GNN are shown in Appendix C.
@@
$\mathcal{R}_{i}$ $\sqrt{\eta^{2}+\phi^{2}}$
tracker type edge between node $i$ and $j$ $|r_{i}-r_{j}|$, $r_{i}r_{j}$, $\eta_{i}-\eta_{j}$, $\phi_{i}-\phi_{j}$, $\mathcal{R}_{i}-\mathcal{R}_{j}$
@@
4 Systematic Uncertainties
The dominant uncertainty source comes from the training variability of neural networks. To assess the robustness of our machine-learning models, 50 CNNs are trained with different initial seeds. The training uncertainty for the neural networks is estimated as half of the difference between the maximum and minimum efficiencies observed, which amounts to approximately $1.7\%$ 11 A similar procedure was applied to the GNN models, yielding a training uncertainty of approximately 2.3%, although this is not included in the final results since the CNN model serves as the baseline model.
. The uncertainty originating from the integrated luminosity for CEPC at $\sqrt{s}=240$ GeV is estimated to be $0.13\%$ [49], which is negligible in this study.
5 Results
Both CNN and GNN methods have been evaluated for the efficiency of LLP signals (type I and II signals combined) from the XGboost output score while keeping background free, the efficiency results are summarized in Table 4. The signal efficiencies of the CNN method range from 30% to 95%, while the GNN method efficiencies range from 29% to 92% for various LLP masses and lifetimes. The signal efficiency is multiplied by the signal acceptance to obtain the signal yield, as shown in Figure 6.
Table 4: Signal efficiencies for different LLPs masses and lifetimes obtained with CNN-based and GNN-based methods. All efficiency values are presented with their corresponding statistical uncertainties.
@@ 5 Results
Figure 6: Efficiency and acceptance of LLP Detection with CEPC detector. The product of geometry acceptance and selection efficiency for LLPs with varying masses and lifetimes. Solid symbols represent data from a CNN-based approach, while hollow symbols are from a GNN-based approach. The varied shapes correspond to different assumed LLP masses, with circles for 50 GeV, diamonds for 10 GeV, and crosses for 1 GeV. The error bars representing statistical uncertainties have been omitted from the figure as their values are too small for effective visual representation.
The signal efficiencies for both the CNN and GNN approaches were evaluated. Their performance is comparable for low-mass LLP scenarios; however, at intermediate and higher LLP masses, the CNN consistently outperforms the GNN. Consequently, CNN is chosen as the baseline model for the subsequent sensitivity analysis and exclusion limit calculations, as reflected in Table 4.
To estimate the sensitivity, we generate 10,000 pseudo-experiments under the null hypothesis (no LLP signal), using the Asimov dataset. In each pseudo-experiment, event yields are randomly sampled from Poisson distributions with means equal to the expected Standard Model background yields, thereby realistically modeling statistical fluctuations. The profile likelihood ratio is employed as the test statistic, defined as: $q_{\mu}=-2\ln\frac{\mathcal{L}(\mu,\hat{\hat{\theta}})}{\mathcal{L}(\hat{\mu},\hat{\theta})},$ where $\mu$ is the signal strength under test, $\hat{\mu}$ and $\hat{\theta}$ denote the best-fit signal strength and nuisance parameters, and $\hat{\hat{\theta}}$ represents the nuisance parameters’ conditional maximum likelihood estimates for a given $\mu$.
The 95% confidence level upper limits on the branching ratio $\mathcal{B}(H\rightarrow XX)$ are derived using the CLs method [50], comparing the observed test statistics against the expected distributions under the null and alternative (with LLP signal) hypotheses. The analyzed samples correspond to a luminosity of $20~{}\text{ab}^{-1}$, corresponding to approximately $4\times 10^{6}$ Higgs bosons [51]. We consider two LLP signal scenarios:
@@
  • Type I and Type II signal yields have a fixed ratio defined by the parameter $\epsilon_{V}:=\frac{BR(X\to\nu\bar{\nu})}{BR(X\to q\bar{q})}$, which is set to 0.2. A one-dimensional 95% confidence level upper limit on $\mathcal{B}(H\rightarrow XX)$ is derived and shown in Figure 7(a). Further detailed results from the pseudo-experiments under the null hypothesis are provided in Appendix D.
@@
  • Type I and Type II signal yields have a floating ratio $\epsilon_{V}$ with an allowed range from $10^{-7}$ to $1$. A one-dimensional 95% Confidence Level upper limit on $\mathcal{B}(H\rightarrow XX)$ is derived and shown in Figure 7b).
@@
  • The ratio of Type I to Type II signal yields is floating, and two-dimensional 95% Confidence Level upper limits on $\mathcal{B}_{\textrm{2-jet}}$ and $\mathcal{B}_{\textrm{4-jet}}$ are derived. A bivariate statistical fit is performed to derive the upper limits and results are shown in Figure 8. More detailed results on 2-D upper limits can be found in Appendix E.
@@
Both 1-D and 2-D exclusion limits are summarized in Table 5. As the results indicate, statistical uncertainties dominate in this analysis.
In addition, we relaxed the background-free assumption to evaluate the lower bound of the experimental sensitivity. The results under the assumption of two background events are presented in Appendix F.
@@
Figure 7: One-dimensional constraints on Higgs boson decays to LLPs. 95% C.L. upper limit on the branching ratio (BR) for the Higgs boson ($H$) decay into pairs of LLPs ($XX$), where $\epsilon_{V}$ is the ratio $\frac{BR(X\to\nu\bar{\nu})}{BR(X\to q\bar{q})}$. a): a fixed ratio $\epsilon_{V}=0.2$, b): a floating $\epsilon_{V}$. The shaded areas indicate statistical and systematic uncertainties combined.
@@
Figure 8: Two-dimensional constraints on Higgs Boson decays to LLPs. The 95% C.L. 2-D upper limit on ($\mathcal{B}_{\textrm{2-jet}}$, $\mathcal{B}_{\textrm{4-jet}}$) for three LLPs masses 50 GeV (left), 10 GeV (middle), 1 GeV (right). Different colored lines indicate different LLP lifetimes. The uncertainties on the limits are omitted and a few limits are scaled by a factor for better visibility: 1) 4 for $M_{X}=10$, $\tau=100$; 2) 0.2 for $M_{X}=1$, $\tau=10$; 3) 4 for $M_{X}=50$, $\tau=100$.
@@
Table 5: The 95% C.L. exclusion limit on BR($h\to XX$) for all signal channels with both fixed and floating $\epsilon_{V}$. The limits include ±$\pm 1\sigma$ uncertainties after taking into account both statistical and systematic contributions.
@@
Fixed 1 $2.72^{+0.41}_{-0.27}$ $1.72^{+0.31}_{-0.15}$ $4.76^{+0.58}_{-0.53}$ $25.05^{+2.12}_{-1.44}$ $1142.42^{+159.89}_{-69.39}$
10 $3.27^{+0.68}_{-0.29}$ $1.38^{+0.19}_{-0.14}$ $1.55^{+0.26}_{-0.16}$ $3.31^{+0.56}_{-0.33}$ $22.87^{+1.85}_{-2.03}$
50 $1.39^{+0.21}_{-0.17}$ $1.02^{+0.17}_{-0.08}$ $1.04^{+0.16}_{-0.09}$ $1.04^{+0.16}_{-0.09}$ $2.76^{+0.45}_{-0.25}$
Floating 1 $2.62^{+0.62}_{-0.35}$ $1.60^{+0.61}_{-0.10}$ $4.26^{+0.57}_{-0.39}$ $23.14^{+4.18}_{-1.16}$ $1030.25^{+153.48}_{-29.36}$
10 $3.14^{+0.92}_{-0.27}$ $1.31^{+0.38}_{-0.12}$ $1.48^{+0.22}_{-0.21}$ $3.16^{+0.47}_{-0.44}$ $20.36^{+3.04}_{-0.79}$
50 $1.31^{+0.32}_{-0.13}$ $1.00^{+0.30}_{-0.09}$ $0.97^{+0.28}_{-0.08}$ $0.98^{+0.28}_{-0.08}$ $2.53^{+0.37}_{-0.26}$
6 Discussion
6.1 Comparison between CNN-based and GNN-based Methods and Selection-based Method
From the signal efficiencies listed in Table 4, both the CNN and GNN approaches deliver excellent performance across the parameter space. Notably, their efficiencies are comparable at low LLP masses; however, as the mass increases, the CNN consistently outperforms the GNN. This mass-dependent difference likely stems from CNNs’ ability to effectively capture spatially localized features in the two-dimensional image representation, which become more pronounced at higher masses. Consequently, we adopt CNN as the baseline for our subsequent sensitivity evaluation and exclusion limit calculations, while acknowledging that GNN remains a competitive alternative—particularly in low-mass regimes. Both methods demonstrate superior capabilities in deciphering the complex data distributions present in the detector space, capturing intricate event structures and topologies more effectively when comparing with the traditional selection-based approach. Detailed comparisons with the selection-based method is available in Appendix G.
CNN is adept at recognizing local features in two-dimensional space, whereas GNN delves into the relationships and structural nuances of nodes within a higher-dimensional framework, effectively identifying displaced vertices, a key characteristic of LLPs. In practical application, CNNs process events as two-dimensional images, while GNNs require an initial transformation of event information into graph structures. This preprocessing step can affect GNN’s performance, contingent on the algorithm’s ability to retain crucial event information. In terms of training speed, GNNs generally achieve faster speeds due to the reduced number of nodes compared to the number of pixels in CNN images. Nevertheless, CNNs tend to achieve quicker convergence, usually within five epochs, as opposed to GNN’s requirement of 15-20 epochs.
To further enhance network performance, GNN could benefit from refined algorithms that improve graph conversion and granularity, all the while preserving data integrity and minimizing graph complexity. For CNN, the accuracy of the network is significantly influenced by image resolution, highlighting computational power as a primary challenge.
6.2 Comparisons with other searches
We evaluate the effectiveness of our ML-based approach for LLPs detection at future lepton colliders by comparing our results with results at hadron colliders with the ATLAS experiment [15, 14] and CMS experiment [16, 17] as well as future HL-LHC experiments [52]. The comparison is based on four primary metrics: signal acceptance, selection efficiency, analysis strategy and signal yields:
@@ 6.2 Comparisons with other searches
  • Selection Efficiency: For hadron colliders, LLPs events typically trigger on displaced decays and/or large missing transverse energy. The LLPs trigger efficiency at the ATLAS experiment is estimated to be between $10^{-3}$ and $0.3$ [15]. Besides trigger efficiency, there are additional efficiencies involved such as displaced vertex/object reconstruction efficiencies which are typically in the order of a few percent. In contrast, LLPs event selection at lepton colliders can adapt to a triggerless-equivalent approach [53, 54] owing to the clean environment. The ML-based approach can be applied directly with low-level detector information without any event-level reconstruction. As a result, our ML-based approach with lepton colliders can achieve an overall selection efficiency as high as $95\%$, an improvement of several orders in magnitude when compared with LHC or HL-LHC efficiencies.
@@
  • Analysis Strategy: Traditionally, analyses of LLPs conducted elsewhere have employed a selection-based method, which involves categorizing events into multiple subsets with different decay modes and orthogonal signal types. These analyses necessitate manual re-tuning and re-optimization for each subset and different LLPs mass and lifetime configurations. In contrast, deep neural networks can be retrained in automation with a similar setup for each LLPs mass and lifetime, resulting in a simplified analysis framework and higher efficiencies compared to the selection-based method. Additional comparisons can also be found in Appendix G.
@@
  • Signal Yields and Upper Limits Comparison: LHC and HL-LHC can produce a significantly larger number of Higgs bosons compared to lepton colliders. Despite this, higher signal acceptance and selection efficiencies in our ML-based approach compensate for the relatively low number of Higgs bosons. We achieve upper limits as low as $9.7\times 10^{-7}$ on $\mathcal{B}(H\rightarrow XX)$ with $4\times 10^{6}$ Higgs bosons. This upper limit is approximately three orders of magnitude better than the $10^{-3}$ limit observed at the ATLAS and CMS experiments with $10^{7}$ Higgs bosons and it is comparable to the projected HL-LHC limit with about $10^{8}$ expected Higgs bosons.
@@
Besides comparing with LLPs searches at hadron colliders, we have also compared our result with a preliminary LLPs study [55] on the ILC [28] sensitivity with a traditional selection-based method. The ILC sensitivity study searches for long-lived dark photons produced in Higgstrahlung events via the Higgs portal. We have compared our result with the hadronic decay dark photo result since the event signature is similar. We have seen that the signal acceptance factors are similar between ILC and CEPC detectors but the signal efficiencies differ significantly. The signal efficiencies in the ILC study range from $0.1\%$ to $10\%$, which is at least an order of magnitude lower than ours. The upper limits on $\mathcal{B}(H\rightarrow XX)$ in the ILC study are derived under the assumption of $100\%$ truth-level signal efficiency. Under this assumption, two results show similar sensitivities only in the low lifetime region ($<1~{}ns$) of LLPs. In the long lifetime region, for example, for a 1 GeV LLP with a lifetime of a few nanoseconds, our result yields an upper limit of about $5\times 10^{-6}$ which is an order of magnitude better than the ILC’s upper limit of $10^{-4}$.
6.3 Comparison with an External Detector
Finally, we quantify the sensitivity gain from deploying an additional external detector alongside the primary detector. As shown in Appendix H, this configuration can improve the LLP detection reach by up to a factor of 13.7 in certain mass–lifetime regimes, reinforcing the potential value of such extensions.
7 Conclusion
In summary, we have investigated the search for LLPs at future lepton colliders using machine learning techniques with full simulation data samples. We have demonstrated that the efficiency of LLP signal detection can reach up to 95% for LLPs with a mass around 50 GeV and a lifetime of approximately 1 nanosecond, surpassing traditional selection-based methods on almost all fronts. Our methodology’s effectiveness is not confined to theoretical constructs but is evidenced by detailed simulations. With a luminosity of $20~{}\text{ab}^{-1}$ and around $4\times 10^{6}$ Higgs bosons analyzed, this study sets new benchmarks for sensitivity in LLP searches with lepton colliders, the branching ratio of Higgs decaying into LLPs reaches $9.7\times 10^{-7}$.
This work not only demonstrates the substantial potential of deep learning techniques in particle physics research but also sets a solid foundation for future explorations at lepton colliders such as CEPC, ILC, FCC, and CLIC. The adaptability of our machine learning approach across different collider environments augurs well for its application in broader physics analyses, promising a new era of discoveries.